OpenStudy (anonymous):

what is the inverse of f if f(x)=^3 sqrt x-4

OpenStudy (anonymous):

i cant help u because u block me

OpenStudy (kc_kennylau):

Do you mean \(\Large f(x)=x^3\sqrt{x-4}\)?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

Do you know what the first step is?

OpenStudy (anonymous):

good night

OpenStudy (anonymous):

Ranado, you kept asking me if I had a boyfriend, when I already gave you my answer. You were annoying me.

OpenStudy (anonymous):

@kc_kennylau No. It's just exactly how I wrote it.

OpenStudy (kc_kennylau):

OHHHHHHHHHHH

OpenStudy (kc_kennylau):

Do you mean \(\Large f(x)=\sqrt[3]{x-4}\)?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

Do you know the first step?

OpenStudy (anonymous):